# Structural properties of nested set complexes

### v1

@@ Abstract.
We study structural and topological properties of nested set complexes of matroids with arbitrary building sets, proving that these complexes are vertex decomposable and admit convex ear decompositions. These results unify and generalize several recent and classical theorems on Bergman complexes and augmented Bergman complexes of matroids. As a first application, we show that the $h$-vector of a nested set complex is strongly flawless and, in particular, top-heavy. We then specialize to the boundary complex of the Deligne–Mumford–Knudsen moduli space $\overline{\mathcal{M}}_{0,n}$ of rational stable marked curves, which coincides with the complex of trees, establishing new structural decomposition theorems and deriving combinatorial formulas for its face enumeration polynomials.
@@ 1.2. Main results
The property of being vertex decomposable, as well as the property of being shellable, is not topological. In particular, we cannot rely on a topological reasoning like the one discussed above to establish the Cohen–Macaulayness of $\mathcal{N}(\mathcal{L},\mathcal{G})$. By providing a careful combinatorial analysis of the behavior of links and deletions of nested set complexes, we establish the following main result.
@@
The inequalities in (1) follow from results by Chari [Cha97, Corollary 2]. The fact that the $g$-vector of a simplicial complex with a convex ear decomposition is an $M$-vector was proved by Swartz in [Swa06, Corollary 3.10]. The fact that the complementary vector of a simplicial complex with a convex ear decomposition is a sum of $M$-vectors is also due to Swartz [Swa06, Proposition 4.4].
@@ Acknowledgments
We thank Spencer Backman, Galen Dorpalen-Barry, Anastasia Nathanson, Ethan Partida, and Noah Prime for kindly discussing their results on shellability in [BDBN+26]. We are also grateful to Christos Athanasiadis for valuable discussions on convex ear decompositions.
@@ 2.3. Convex ear decomposition
Swartz [Swa06, Theorem 4.1] proved that every simplicial complex $\Delta$ with a convex ear decomposition is doubly Cohen–Macaulay, over any field of characteristic 0. Despite the fact that double Cohen–Macaulayness is a topological property, convex ear decompositions are easily seen to not be preserved under homeomorphisms.
@@
$\Delta_{1}$$\Delta_{1}\cup\Delta_{2}$$\Delta_{2}$$\Delta_{1}\cup\Delta_{2}\cup\Delta_{3}$$\Delta_{3}$
Figure 3. Attaching two 22-dimensional simplicial balls, or “ears” to a simplicial polytope $\Delta_{1}$.
@@ 2.4. Face enumeration
The $g$-vector of $\Delta$, denoted $(g_{0}(\Delta),\ldots,g_{\lfloor d/2\rfloor}(\Delta))$ is defined by $g_{0}(\Delta)=h_{0}(\Delta)$, and $g_{i}(\Delta)=h_{i}(\Delta)-h_{i-1}(\Delta)$ for $1\leq i\leq\lfloor d/2\rfloor$. The complementary vector of $\Delta$, which we denote by $(c_{0}(\Delta),\ldots,c_{\lfloor d/2\rfloor}(\Delta))$ is defined by $c_{i}(\Delta)=h_{d-i}(\Delta)-h_{i}(\Delta)$ for each $0\leq i\leq\lfloor d/2\rfloor$.
@@
By using ideas related to the $g$-conjecture for simplicial spheres, Adiprasito, Papadakis, and Petrotou announced a result [APP21, Corollary 3.2] that implies that the first part of the above theorem continues to hold for all doubly Cohen–Macaulay complexes. On the other hand, a recent paper by Larson and Stapledon implies the second part of the above statement for doubly Cohen–Macaulay complexes, and additionally provides a characterization of all complementary vectors of simplicial complexes admitting a convex ear decomposition [LS25, Corollary 1.7].
@@ 5. Braid matroids and the minimal building set
Let $\Delta$ be the nested set complex $\mathcal{N}(\Pi_{4},\mathcal{G}_{\min})$ which is the $1$-dimensional simplicial complex of the Petersen graph, and $\mathrm{cone}(\Delta)$ its cone of pure dimension $2$; see Figure˜6. By counting faces, the $f$-vector of $\Delta$ is $f(\Delta)=(1,10,15)$. Using
@@
For every $0\leqslant k\leqslant n$, let us denote by $Q_{n,k}$ the number of Stirling permutations $\pi\in Q_{n}$ that have exactly $k$ descents. Then, $Q_{n,k}\leqslant Q_{n,n-k}$ for each $k\leqslant\lfloor n/2\rfloor$ and $Q_{n,0}\leqslant Q_{n,1}\leqslant\cdots\leqslant Q_{n,\lfloor n/2\rfloor}$.
It is natural to ask if the polynomials $P_{n}(x)$ are unimodal for each $n$. The answer is affirmative and, in fact, much more is true. Thanks to a result of Bóna [Bón09], these polynomials are real-rooted. Real-rootedness implies (ultra) log-concavity without internal zeros, which in turn implies the unimodality of these polynomials (see [Sta89, Brä15] for more on these classes of inequalities). One of the consequences of the above corollary is that the mode is located on the “top half” of the polynomials.
@@ 6.2. Two problems on Stirling permutations
Find suitable injections $Q_{n,k}\hookrightarrow Q_{n,n-k}$ for $k\leq\lfloor n/2\rfloor$ and $Q_{n,k}\hookrightarrow Q_{n,k+1}$ for $0\leq k\leq\lfloor n/2\rfloor-1$.
@@
6.3. Convex ear decompositions and shellability
@@
For any geometric lattice $\mathcal{L}$ and any building set $\mathcal{G}$, does the nested set complex $\mathcal{N}(\mathcal{L},\mathcal{G})$ admit a PS ear decomposition?
@@
Whilel we are inclined to expect such examples to abound, we have not been able to construct one.

### v2

@@ Abstract.
We study structural and topological properties of nested set complexes of matroids with arbitrary building sets, proving that these complexes are vertex decomposable and admit convex ear decompositions. These results unify and generalize several recent and classical theorems about nested set complexes, Bergman complexes, and augmented Bergman complexes of matroids. As a first application, we show that the $h$-vector of a nested set complex is strongly flawless and, in particular, top-heavy. We then specialize to the boundary complex of the Deligne–Mumford–Knudsen moduli space $\overline{\mathcal{M}}_{0,n}$ of rational stable marked curves, which coincides with the complex of trees, establishing new structural decomposition theorems and deriving combinatorial formulas for its face enumeration polynomials.
@@ 1.2. Main results
The property of being vertex decomposable, as well as the property of being shellable, is not closed under homeomorphisms. In particular, we are not able to use a topological reasoning like the one used to establish the Cohen–Macaulayness of $\mathcal{N}(\mathcal{L},\mathcal{G})$ as above. By providing a careful combinatorial analysis of the behavior of links and deletions of nested set complexes, we establish the following main result.
@@
The inequalities in (1) follow from results by Chari [Cha97, Corollary 2]. The fact that the $g$-vector of a simplicial complex with a convex ear decomposition is an $M$-vector was proved by Swartz in [Swa06, Corollary 3.10]. The fact that the complementary vector of a simplicial complex with a convex ear decomposition is a sum of $M$-vectors follows from results by Swartz [Swa06, Section 4] or, in a more general context, by work of Larson and Stapledon [LS25].
@@ Acknowledgments
We thank Spencer Backman, Galen Dorpalen-Barry, Anastasia Nathanson, Ethan Partida, and Noah Prime for kindly discussing their results on shellability in [BDBN+26]. We are also grateful to Christos Athanasiadis for valuable discussions on convex ear decompositions, and to Matt Larson for comments on doubly Cohen–Macaulay complexes.
@@ 2.3. Convex ear decomposition
Figure 3. A convex ear decomposition of an “ice cream cone,” consisting of $\Delta_{1}$ (the boundary of the octahedron), and the “ears” $\Delta_{2}$ and $\Delta_{3}$, which are 22-dimensional simplicial balls.
Swartz [Swa06, Theorem 4.1] proved that every simplicial complex $\Delta$ with a convex ear decomposition is doubly Cohen–Macaulay over $\mathbb{Z}$. Despite the fact that double Cohen–Macaulayness is a topological property, convex ear decompositions are easily seen to not be preserved under homeomorphisms.
@@ 2.4. Face enumeration
The $g$-vector of $\Delta$, denoted $(g_{0}(\Delta),\ldots,g_{\lfloor d/2\rfloor}(\Delta))$ is defined by $g_{0}(\Delta)=h_{0}(\Delta)$, and $g_{i}(\Delta)=h_{i}(\Delta)-h_{i-1}(\Delta)$ for $1\leqslant i\leqslant\lfloor d/2\rfloor$. The complementary vector of $\Delta$, which we denote by $(c_{0}(\Delta),\ldots,c_{\lfloor d/2\rfloor}(\Delta))$ is defined by $c_{i}(\Delta)=h_{d-i}(\Delta)-h_{i}(\Delta)$ for each $0\leqslant i\leqslant\lfloor d/2\rfloor$.
@@
By using ideas related to the $g$-conjecture for simplicial spheres, Adiprasito, Papadakis, and Petrotou announced a result [APP21, Corollary 3.2] that implies that the first part of the above theorem continues to hold for doubly Cohen–Macaulay complexes. A different proof is provided in a recent paper by Larson and Stapledon, which in addition implies a stronger version of the second part of the above statement for doubly Cohen–Macaulay complexes over a field of characteristic $2$ [LS25, Corollary 1.7]. Furthermore, Larson and Stapledon provide a characterization of all complementary vectors of simplicial complexes admitting a convex ear decomposition, and deduce additional inequalities for $h$-vectors of doubly Cohen–Macaulay complexes over a field of characteristic $2$ [LS25, Corollary 4.5].
@@ 4. Convex ear decomposition
Figure 6. A convex ear decomposition of $\mathcal{N}(\mathrm{U}_{3,4},\mathcal{G}_{\max})$, for the uniform matroid on of rank 33 on 44 elements, with the maximal building set.
@@ 5. Braid matroids and the minimal building set
Let $\Delta$ be the nested set complex $\mathcal{N}(\Pi_{4},\mathcal{G}_{\min})$ which is the $1$-dimensional simplicial complex of the Petersen graph, and $\mathrm{cone}(\Delta)$ its cone of pure dimension $2$; see Figure˜7. By counting faces, the $f$-vector of $\Delta$ is $f(\Delta)=(1,10,15)$. Using
@@
For every $0\leqslant k\leqslant n$, let us denote by $Q_{n,k}$ the set of Stirling permutations $\pi\in Q_{n}$ that have exactly $k$ descents. Then, $|Q_{n,k}|\leqslant|Q_{n,n-k}|$ for each $k\leqslant\lfloor n/2\rfloor$ and $|Q_{n,0}|\leqslant|Q_{n,1}|\leqslant\cdots\leqslant|Q_{n,\lfloor n/2\rfloor}|$.
It is natural to ask if the polynomials $P_{n}(x)$ are unimodal for each $n$. The answer is affirmative and, in fact, much more is true. Thanks to a result of Bóna [Bón09], these polynomials are real-rooted. Real-rootedness implies (ultra) log-concavity without internal zeros, which in turn implies the unimodality of these polynomials (see [Sta89, Brä15] for more on these classes of inequalities). One of the consequences of the above corollary is that the mode is located at the “top half” of the polynomials.
@@ 6.2. Two problems on Stirling permutations
Find suitable injections $Q_{n,k}\hookrightarrow Q_{n,n-k}$ for $k\leqslant\lfloor n/2\rfloor$ and $Q_{n,k}\hookrightarrow Q_{n,k+1}$ for $0\leqslant k\leqslant\lfloor n/2\rfloor-1$.
@@
6.3. Convex ear decompositions, and shellability
@@
For which geometric lattices $\mathcal{L}$ and which building sets $\mathcal{G}$, does the nested set complex $\mathcal{N}(\mathcal{L},\mathcal{G})$ admit a PS ear decomposition?
In general, $\mathcal{N}(\mathcal{L},\mathcal{G})$ does not admit a PS ear decomposition. For example, the order complex of $\mathcal{L}(U_{3,4})$ only admits the convex ear decomposition shown in Figure˜6 up to isomorphism, which is not a PS ear decomposition.
@@
While we are inclined to expect such examples to abound, we have not been able to construct one.